OpenStudy (crashonce):

Find the least integral value of n for the following product to exceed 10^5

OpenStudy (crashonce):

\[10^{\frac{ 1 }{ 11 }} \times 10^{\frac{ 2 }{ 11 }} \times 10^{\frac{ 3 }{ 11 }}.......10^{\frac{ n }{11 }}\]

OpenStudy (crashonce):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

hmm, well you can add the exponents to get \[\LARGE 10^{\frac{1}{11}+\frac{2}{11}+\cdots\frac{n}{11}}\] \[\LARGE 10^{\frac{1+2+\cdots n}{11}}\] \[\LARGE 10^{\frac{\frac{n(n+1)}{2}}{11}}\] \[\LARGE 10^{\frac{n(n+1)}{22}}\]

jimthompson5910 (jim_thompson5910):

you don't want to exceed 10^5, so the exponent is at most 5 \[\Large \frac{n(n+1)}{22} \le 5\]

jimthompson5910 (jim_thompson5910):

solve for n

OpenStudy (crashonce):

10 right?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (crashonce):

ok thanks

jimthompson5910 (jim_thompson5910):

oh sorry you do want to exceed 10^5 n = 10 will make n(n+1)/22 equal to 5 so n > 10 will make n(n+1)/22 more than 5 the next whole number up is 11, so the answer is n = 11

OpenStudy (crashonce):

ohhhhhh oh ritght yea